# Towards a Cascaded LLM Framework for Cost-effective Human-AI Decision-Making

### v2

@@ 5.2 Online Improvement of the Decision System
(c) MedQA
(d) MedMCQA
Figure 4: Cumulative Regret in Online Setting (Llama 1B–8B). Cumulative system risk over time. Training data is collected only when abstentions occur. The cascaded system consistently achieves lower regret. The shaded area indicates the standard error.
Desideratum D3 requires that “the framework should continuously refine its deferral and abstention policies based on feedback, ensuring risk-minimised decision-making”. We simulate an online setting in which the system selects among $\mathcal{M}_{\text{base}}$, $\mathcal{M}_{\text{large}}$, or a human expert, adjusting its thresholds based on feedback from abstentions.
To avoid trivial solutions (e.g., always selecting one model), we balance system risk using $\lambda_{c}=10^{-5}$ and $\lambda_{a}=0.1$, in line with Zellinger et al. (2025). Feedback is collected in batches and stored in a training buffer. Teh deferral and abstention thresholds $\boldsymbol{\theta}=\{\phi_{\text{base}},\xi_{\text{base}},\xi_{\text{large}}\}$ are updated with the ADAM optimiser (Kingma and Ba, 2014), using a learning rate of 0.05 and batch size 10. Thresholds are initialised at $\boldsymbol{\theta}^{(0)}=\{0.5,0.05,0.05\}$, where $\xi_{i}=0.05$ corresponds to the standard deviation of 5% confidence. For the single model baselines, we initialise the abstention threshold als with $\xi=0.05$, and keep the rest of the hyperparameters the same. Throughout this experiment, we use the STP ($n=1$) verification strategy, which was most competitive in the previous section.
We compare the cascaded system $C$ to using only $\mathcal{M}_{\text{base}}$ and only $\mathcal{M}_{\text{large}}$ with a single abstention threshold $\xi$. The system risk for a single model is explained in more detail in Appendix A.1. As we are considering a deployed system, we track the cumulative regret over time, which we define as follows:
@@
where $\mathcal{M}\in\{C,\mathcal{M}_{\text{base}},\mathcal{M}_{\text{large}}\}$ and $\mathcal{M}^{(t)}$ evolves based on abstention feedback $\mathcal{D}_{t}$.
Figure 4 shows that the cascaded system yields lower cumulative regret over 1000 test samples on ARC2-Easy, ARC2-Challenge, and MedMCQA, compared to using a either model in isolation. On MedQA, gains are less clear, likely due to poor confidence estimation, which were also observed in the section above. Similar trends are observed in other cascades (see Figures  8, 11, 14 in Appendix C). Nevertheless, in three of the four cases, the cascaded LLM system demonstrates lower cumulative regret than when using single models online, where feedback is received when abstained.
@@ 6 Limitations
Our cascaded multi-LLM decision-making framework balances accuracy, cost, and abstention but has limitations. Sensitivity to cost and abstention variations can impact efficiency, leading to trivial solutions (only using the cheapest model or the model with the lowest error rate). Discrepancies in model performance or relative costs may cause over-reliance on specific models, reducing adaptability. Furthermore, parameter initialisation affects deferral policy convergence. Additionally, the framework depends on human feedback, which may hinder adaptation if sparse or noisy in a real-world scenario. Finally, fitting a Bayesian logistic regression is usually slightly more complex than a regular one, depending on different posterior approximations or sampling strategies.
@@ 7 Conclusion
We proposed a multi-tier decision-making framework that escalates tasks between a base model, a large model, and human experts. By leveraging deferral and abstention policies, our approach aims to enhance performance, accuracy, and abstention while adapting through online learning. Our experiments show that the framework outperforms single-model baselines by reducing unnecessary escalations and improving response correctness on the ARC2-Easy, ARC2-Challenge, and MedMCQA datasets. On MedQA a cascaded model did not outperform the single model approach, potentially due to the complexity of the dataset. Nevertheless, we believe that this proposed system could be beneficial where performance, costs, and abstention of LLMs need to be carefully balanced. Potential applications for this system are in high-stakes environments such as medicine, trading, and education. In this work, we focused on different methods to calculate confidence scores of correct answers, using the posterior predictive’s standard deviation as uncertainty. Future work should investigate different uncertainty quantification methods of LLMs to enhance abstention. Moreover, it would be crucial to examine if there are theoretical guarantees that justify applying cascaded LLMs.
Acknowledgements
We would like to extend our gratitude to Yusuke Kano, Jeremy Voisey, and Alison O’Neil for their insightful discussion and valuable feedback. Claudio Fanconi’s studentship is funded by Canon Medical Systems Corporation.
@@ Appendix B Implementation Details
B.1 Prompts
@@ Appendix C Additional Results
C.1 LLama 1B $\rightarrow$ 8B
@@
SV ($n{=}5$) 143.4 $\pm$ 33.6 50.1 $\pm$ 13.7 83.1 $\pm$ 31.6 10.7 $\pm$ 11.8 10.1 $\pm$ 10.2 -8.7 $\pm$ 9.1 63.5 $\pm$ 32.0 -3.1 $\pm$ 9.0
STP ($n{=}1$) 129.4 $\pm$ 19.1 118.3 $\pm$ 15.8 59.1 $\pm$ 16.5 38.1 $\pm$ 13.7 7.3 $\pm$ 9.9 5.6 $\pm$ 9.7 30.0 $\pm$ 11.8 -3.7 $\pm$ 9.2
MC-STP ($n{=}5$) 71.0 $\pm$ 17.2 97.7 $\pm$ 14.4 24.0 $\pm$ 16.5 15.4 $\pm$ 11.7 4.2 $\pm$ 10.0 -6.8 $\pm$ 9.2 9.2 $\pm$ 13.2 2.6 $\pm$ 9.0
Table 2: $\Delta$IBC scores for LLaMA (1B–8B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. Uncalibrated values are preserved from the original, calibrated values are updated. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
@@
C.2 Qwen 1.5B $\rightarrow$ 7B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 278.5 $\pm$ 237.3 -2.4 $\pm$ 32.6 80.5 $\pm$ 174.1 -11.6 $\pm$ 20.5 -45.8 $\pm$ 118.8 -27.3 $\pm$ 14.3 70.4 $\pm$ 89.0 -24.8 $\pm$ 9.5
SV ($n{=}5$) 93.3 $\pm$ 112.7 10.2 $\pm$ 29.6 -15.6 $\pm$ 104.2 -26.2 $\pm$ 19.1 -65.6 $\pm$ 55.3 -25.4 $\pm$ 13.4 29.3 $\pm$ 63.2 -19.5 $\pm$ 9.4
STP ($n{=}1$) -100.0 $\pm$ 948.0 -16.9 $\pm$ 33.2 -100.0 $\pm$ 847.4 -19.3 $\pm$ 20.9 -87.0 $\pm$ 313.1 -23.9 $\pm$ 14.7 -100.0 $\pm$ 835.5 -22.9 $\pm$ 9.7
MC-STP ($n{=}5$) -66.3 $\pm$ 142.6 -37.4 $\pm$ 28.8 -74.3 $\pm$ 143.4 -16.0 $\pm$ 21.0 -69.7 $\pm$ 72.9 -27.3 $\pm$ 13.7 -71.6 $\pm$ 114.7 -22.2 $\pm$ 9.4
[…]
SV ($n{=}5$) 219.7 $\pm$ 35.8 188.5 $\pm$ 31.5 66.5 $\pm$ 22.3 51.0 $\pm$ 23.7 13.9 $\pm$ 18.9 -11.9 $\pm$ 14.7 24.6 $\pm$ 13.8 -11.0 $\pm$ 9.8
STP ($n{=}1$) 284.5 $\pm$ 40.4 242.7 $\pm$ 36.3 96.9 $\pm$ 24.2 89.3 $\pm$ 26.2 21.6 $\pm$ 19.4 1.7 $\pm$ 16.3 35.0 $\pm$ 14.2 -10.0 $\pm$ 10.1
MC-STP ($n{=}5$) 66.5 $\pm$ 19.2 171.4 $\pm$ 36.4 21.3 $\pm$ 14.6 39.2 $\pm$ 22.0 -10.5 $\pm$ 14.6 -16.8 $\pm$ 14.5 -3.5 $\pm$ 10.2 -19.7 $\pm$ 9.3
Table 3: $\Delta$IBC scores for Qwen (1.5B–7B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. Uncalibrated values are preserved from the original, calibrated values are updated. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.3 LLama 3B $\rightarrow$ 8B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 267.7 $\pm$ 364.0 3.8 $\pm$ 96.2 322.3 $\pm$ 335.2 -11.5 $\pm$ 81.4 15.7 $\pm$ 205.7 -48.6 $\pm$ 46.9 93.6 $\pm$ 154.5 -27.8 $\pm$ 20.9
SV ($n{=}5$) 247.2 $\pm$ 359.8 32.1 $\pm$ 100.9 352.2 $\pm$ 325.3 0.0 $\pm$ 84.4 12.5 $\pm$ 200.0 -28.2 $\pm$ 49.4 96.1 $\pm$ 146.7 -39.5 $\pm$ 20.4
STP ($n{=}1$) 272.7 $\pm$ 338.6 34.9 $\pm$ 99.6 279.7 $\pm$ 337.5 -40.7 $\pm$ 80.9 -35.6 $\pm$ 199.1 -8.2 $\pm$ 49.3 91.7 $\pm$ 153.0 -19.2 $\pm$ 21.9
MC-STP ($n{=}5$) 272.4 $\pm$ 325.0 44.0 $\pm$ 96.7 279.9 $\pm$ 312.6 33.0 $\pm$ 100.0 -9.2 $\pm$ 204.9 -23.2 $\pm$ 52.8 119.0 $\pm$ 154.3 -45.0 $\pm$ 20.0
[…]
SV ($n{=}5$) 257.4 $\pm$ 324.7 63.2 $\pm$ 106.4 281.6 $\pm$ 292.7 37.4 $\pm$ 83.1 4.7 $\pm$ 131.5 -32.4 $\pm$ 48.8 84.8 $\pm$ 107.0 -14.2 $\pm$ 22.1
STP ($n{=}1$) 331.9 $\pm$ 287.7 77.0 $\pm$ 112.2 254.6 $\pm$ 291.8 -0.7 $\pm$ 83.7 -8.8 $\pm$ 126.8 -34.4 $\pm$ 46.2 95.7 $\pm$ 99.5 -38.6 $\pm$ 20.3
MC-STP ($n{=}5$) 422.2 $\pm$ 328.8 78.2 $\pm$ 109.8 325.5 $\pm$ 306.1 -6.6 $\pm$ 78.7 65.2 $\pm$ 150.7 -52.8 $\pm$ 48.1 91.4 $\pm$ 100.3 -11.5 $\pm$ 22.1
Table 4: $\Delta$IBC scores for LLaMA (3B–8B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA

### v3

@@ 5.2 Online Improvement of the Decision System
Desideratum D3 requires that “the framework should continuously refine its deferral and abstention policies as feedback becomes available, ensuring sustained reliability and improvement”. We simulate an online setting in which the system selects among $\mathcal{M}_{\text{base}}$, $\mathcal{M}_{\text{large}}$, or a human expert, adjusting its thresholds based on feedback from abstentions.
@@
(c) MMLU
(d) MedQA
(e) MedMCQA
Figure 4: Cumulative Regret in Online Setting (Qwen-2.5 1.5B $\rightarrow$ 7B). Cumulative system risk over time. Training data is collected only when abstentions occur. The cascaded system consistently achieves lower regret.
Specifically, the experiment streams 1,000 unseen questions in a random order to simulate production traffic. Before the first query, the deferral probabilities are calibrated using the values learned from the 100-sample calibration set. Thereafter, we add queries that were marked for abstention and answered by an oracle expert to a replay buffer. This replay buffer is used to perform the ADAM optimiser (Kingma and Ba, 2014) updates on the differentiable risk 2 with a learning rate of 0.05 and a batch size of 10, on the deferral and abstention thresholds $\boldsymbol{\theta}=\{\phi_{\text{base}},\xi_{\text{base}},\xi_{\text{large}}\}$. The prediction is made on an unseen query and the regret is calculated on it. If a query is added to the replay buffer, the regret associated with it has already been calculated, and the sample now becomes a training sample; however, it is not further evaluated. We compare the cascaded system $C$ to using only $\mathcal{M}_{\text{base}}$ and only $\mathcal{M}_{\text{large}}$ with a single abstention threshold $\xi$. The system risk for a single model is explained in more detail in Appendix A.1. As we are considering a deployed system, we track the cumulative regret over time, which we define as follows:
@@
where $\mathcal{M}\in\{C,\mathcal{M}_{\text{base}},\mathcal{M}_{\text{large}}\}$ and $\mathcal{M}^{(t)}$ evolves based on abstention feedback $\mathcal{D}_{t}$. We chose regret as the metric for this experiment, inspired by the work on online decision mediation (Jarrett et al., 2022). Regret is the running sum of our per‑query risk. Because error, compute, and human‑hand‑off are already weighted into the same units, adding them over time tells you the exact “bill” the system has paid. A lower regret curve indicates a higher benefit, as it represents a combination of abstentions, correct predictions, and costs, and we can see it grow over time in a deployed setting. The regret curve illustrates how quickly a policy learns online and whether early mistakes are compensated for later.
We initialise the thresholds at $\boldsymbol{\theta}^{(0)}=\{0.5,0.05,0.05\}$, where $\xi_{i}=0.05$ corresponds to the standard deviation of 5% confidence. For the single model baselines, we initialise the abstention threshold as with $\xi=0.05$, and keep the rest of the hyperparameters the same. Throughout this experiment, we employ the STP ($n=1$) verification strategy, which was found to be the most competitive in the previous section. To avoid trivial solutions (e.g., always selecting one model), we balance system risk using $\lambda_{c}=10^{-5}$ and $\lambda_{a}=0.1$, in line with Zellinger et al. (2025).
Figure 4 shows that the cascaded system yields lower cumulative regret over 1000 test samples on ARC2-Easy, ARC2-Challenge, MMLU, and MedMCQA, compared to using either model in isolation. On MedQA, gains are less clear, likely due to poor confidence estimation, which was also observed in the section above. Similar trends are observed in other cascades (see Figures 9, 12, 15 in Appendix C). Nevertheless, in four of the five cases, the cascaded LLM system demonstrates lower cumulative regret than when using single models online, where feedback is received when abstaining from action. Additionally, we experiment by comparing our proposed gradient-based approach to a traditional grid search over $\mathbf{\theta}$ in Appendix C.7.
5.3 The Effect of Imperfect Expert
Figure 5: Imperfect Experts. We increase the percentage of flipped labels during system calibration, simulating imperfect experts, which in turn increases the system’s risk of error.
To recall the system risk objective in Equation 15, the part where wrong human annotations will have an impact is the expected correctness (Equation 13). More precisely, the calibrated confidence scores $\Phi_{\text{base}}$ and $\Phi_{\text{large}}$. The noisier the feedback is, the more uncalibrated $\Phi_{\text{base}}$ and $\Phi_{\text{large}}$ will become. Therefore, optimisation of the cascaded model will become unreliable.
We demonstrate this through an additional experiments in the online setting, where we progressively swap correct to incorrect predictions while calibrating the model, and how this affects the trajectory on the ARC-Easy dataset with Qwen-2.5-1.5B $\rightarrow$ Qwen-2.5-7B. The results are displayed in Figure 5. We observe that the higher the percentage of label corruption is in the calibration set, the higher the cumulative regret becomes while deploying the decision-making system.
@@ 6 Limitations
Our cascaded multi-LLM decision-making framework strikes a balance between accuracy, cost, and abstention, but it has limitations. Sensitivity to cost and abstention variations can impact efficiency, leading to trivial solutions (only using the cheapest model or the model with the lowest error rate). Discrepancies in model performance or relative costs may lead to over-reliance on specific models, thereby reducing adaptability. Furthermore, parameter initialisation affects the convergence of the deferral policy. Additionally, the framework relies on human feedback, which may hinder adaptation if it is sparse or noisy in a real-world scenario. Finally, fitting a Bayesian logistic regression model is usually more complex than fitting a regular one, depending on the different posterior approximations or sampling strategies employed.
@@ 7 Conclusion
We proposed a multi-tier decision-making framework that escalates tasks between a base model, a large model, and human experts. By leveraging deferral and abstention policies, our approach aims to enhance performance, accuracy, and abstention while adapting through online learning. Our experiments show that the framework outperforms single-model baselines by reducing unnecessary escalations and improving response correctness on the ARC2-Easy, ARC2-Challenge, MMLU, and MedMCQA datasets. On MedQA, a cascaded model did not outperform the single model approach, potentially due to the complexity of the dataset. Nevertheless, we believe that this proposed system could be beneficial where performance, costs, and abstention of LLMs need to be carefully balanced. Future work should investigate different uncertainty quantification methods of LLMs to enhance abstention. Moreover, it would be crucial to examine whether there are theoretical guarantees that justify the application of cascaded LLMs.
Acknowledgements and Disclosure of Funding
We want to extend our gratitude to Alan Jeffares, Paulius Rauba, Yusuke Kano, Jeremy Voisey, and Alison Smithard for their insightful discussions and valuable feedback. Canon Medical Systems Corporation funds CF’s studentship. This work was supported by Microsoft’s Accelerate Foundation Models Academic Research initiative.
@@ Appendix B Implementation Details
The code for this paper, to reproduce the results is provided at https://github.com/fanconic/cascaded-llms. All experiments are implemented in Python [Van Rossum and Drake Jr, 1995] with PyTorch [Paszke et al., 2017] and Hugging Face Transformers [Wolf et al., 2020].
Compute.
Experiments are conducted on a single A100-class GPUs.
B.1 Generation Models
Policies are initialised from instruction-tuned checkpoints and trained with the learned reward signal. The following policy backbones are used:
B.2 Prompts
@@
Reasoning:
Response Generation Prompt MMLU
You are an expert in {subject}.
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choices}
@@ Appendix C Additional Results
C.1 Qwen-2.5 1.5B $\rightarrow$ 7B
C.1.1 Uncalibrated
Figure 6: Benefit-Cost Analysis of Uncalibrated Verification Methods (Qwen-2.5 1.5B-7B). We display the cost vs accuracy of the various verification methods, using the cascade ($\texttt{Qwen-2.5-1.5B}\rightarrow\texttt{Qwen-2.5-7B})$. Verification methods, which are located above the linear interpolation between the base or large models, indicate a positive cost-benefit ratio. The error bars indicate the standard error.
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) 278.5 $\pm$ 237.3 80.5 $\pm$ 174.1 -18.7 $\pm$ 4.9 -45.8 $\pm$ 118.8 70.4 $\pm$ 89.0
SV (n=5n{=}5) 93.3 $\pm$ 112.7 -15.6 $\pm$ 104.2 -19.4 $\pm$ 4.9 -65.6 $\pm$ 55.3 29.3 $\pm$ 63.2
STP (n=1n{=}1) -100.0 $\pm$ 948.0 -100.0 $\pm$ 847.4 -20.4 $\pm$ 4.5 -87.0 $\pm$ 313.1 -100.0 $\pm$ 835.5
MC-STP (n=5n{=}5) -66.3 $\pm$ 142.6 -74.3 $\pm$ 143.4 -24.6 $\pm$ 4.2 -69.7 $\pm$ 72.9 -71.6 $\pm$ 114.7
Large SV (n=1n{=}1) 280.5 $\pm$ 41.6 84.8 $\pm$ 24.6 -3.0 $\pm$ 5.0 24.7 $\pm$ 21.0 32.0 $\pm$ 14.0
SV (n=5n{=}5) 219.7 $\pm$ 35.8 66.5 $\pm$ 22.3 -7.0 $\pm$ 4.8 13.9 $\pm$ 18.9 24.6 $\pm$ 13.8
[…]
MC-STP (n=5n{=}5) 66.5 $\pm$ 19.2 21.3 $\pm$ 14.6 -5.7 $\pm$ 4.9 -10.5 $\pm$ 14.6 -3.5 $\pm$ 10.2
[…]
C.2 LLama3 1B $\rightarrow$ 8B
C.2.1 Calibrated
[…]
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) -8.8 $\pm$ 16.4 -13.2 $\pm$ 11.4 -9.6 $\pm$ 2.6 -8.0 $\pm$ 9.5 -9.9 $\pm$ 8.9
SV (n=5n{=}5) -12.8 $\pm$ 16.5 -19.6 $\pm$ 10.8 -11.9 $\pm$ 2.6 -21.3 $\pm$ 8.6 -15.6 $\pm$ 8.7
STP (n=1n{=}1) -4.4 $\pm$ 16.7 -5.2 $\pm$ 11.7 -6.8 $\pm$ 2.8 -13.6 $\pm$ 9.3 -4.8 $\pm$ 9.3
MC-STP (n=5n{=}5) 0.8 $\pm$ 15.2 -9.0 $\pm$ 11.4 -9.0 $\pm$ 2.7 -9.4 $\pm$ 9.2 -11.2 $\pm$ 8.8
Large SV (n=1n{=}1) 38.6 $\pm$ 14.5 2.8 $\pm$ 11.1 -10.6 $\pm$ 2.6 -4.1 $\pm$ 9.9 -0.0 $\pm$ 9.1
SV (n=5n{=}5) 50.1 $\pm$ 13.7 10.7 $\pm$ 11.8 -19.7 $\pm$ 2.3 -8.7 $\pm$ 9.1 -3.1 $\pm$ 9.0
STP (n=1n{=}1) 118.3 $\pm$ 15.8 38.1 $\pm$ 13.7 -1.2 $\pm$ 2.9 5.6 $\pm$ 9.7 -3.7 $\pm$ 9.2
MC-STP (n=5n{=}5) 97.7 $\pm$ 14.4 15.4 $\pm$ 11.7 -11.7 $\pm$ 2.5 -6.8 $\pm$ 9.2 2.6 $\pm$ 9.0
[…]
C.2.2 Uncalibrated
[…]
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) 6.5 $\pm$ 13.1 2.6 $\pm$ 13.8 -9.6 $\pm$ 2.6 -38.9 $\pm$ 98.5 -4.2 $\pm$ 12.0
SV (n=5n{=}5) 16.1 $\pm$ 21.3 -11.3 $\pm$ 18.4 -11.9 $\pm$ 2.6 -55.7 $\pm$ 82.4 -14.4 $\pm$ 18.6
STP (n=1n{=}1) 6.4 $\pm$ 344.3 -30.0 $\pm$ 211.7 -6.8 $\pm$ 2.8 -13.9 $\pm$ 87.0 -21.2 $\pm$ 159.6
MC-STP (n=5n{=}5) -42.5 $\pm$ 116.2 -39.0 $\pm$ 113.6 -9.0 $\pm$ 2.7 -19.8 $\pm$ 107.9 -42.9 $\pm$ 68.9
Large SV (n=1n{=}1) 167.9 $\pm$ 38.0 89.3 $\pm$ 36.6 -7.9 $\pm$ 2.7 9.9 $\pm$ 10.2 73.3 $\pm$ 32.8
SV (n=5n{=}5) 143.4 $\pm$ 33.6 83.1 $\pm$ 31.6 -9.1 $\pm$ 2.6 10.1 $\pm$ 10.2 63.5 $\pm$ 32.0
STP (n=1n{=}1) 129.4 $\pm$ 19.1 59.1 $\pm$ 16.5 1.6 $\pm$ 3.0 7.3 $\pm$ 9.9 30.0 $\pm$ 11.8
MC-STP (n=5n{=}5) 71.0 $\pm$ 17.2 24.0 $\pm$ 16.5 0.3 $\pm$ 2.9 4.2 $\pm$ 10.0 9.2 $\pm$ 13.2
[…]
C.2.3 Online Learning
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MMLU
(d) MedQA
(e) MedMCQA
[…]
C.3 LLama 3B $\rightarrow$ 8B
Due to the large size of the MMLU dataset and the costs associated with making predictions on it, we omit this for the (Llama3.2-3B $\rightarrow$ Llama3.1-8B) combination in this subsection.